Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows: Standard Quantity Standard Frice or Hate $30.00 per ounce $14.00 per hour $3.40 per hour Standard or Hours Direct material:s Direct labor Variable manufacturing overhead Total standard cost per unit 2.00 ounces $60.00 ours 0.50 hours 1. 70 $ 68. 70 During November, the following activity was recorded related to the production of Fludex: a. Materials purchased, 10,000 ounces at a cost of $287,000 b. There was no beginning inventory of materials, however, at the end of the month, 3,000 ounces of material remained in ending c. The company employs 20 lab technicians to work on the production of Fludex. During November, they each worked an average of d. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs e. During November, the company produced 3,400 units of Fludex. inventory. 130 hours at an average pay rate of $12.00 per hour. during November totaled $4,700 Required 1. For direct materials a. Compute the price and quantity variances b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract? 2. For direct labor a. Compute the rate and efficiency variances b. In the past, the 20 technicians employed in the production of Fludex consisted of 4 senior technicians and 16 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued? 3. Compute the variable overhead rate and efficiency variances.
Complete this question by entering your answers in the tabs below. Req 1AReq 1B Req 2AReq 2B Req 3 For direct materials, compute the price and quantity variances. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.) Materials price variance Materials quantity variance Req 1A Req 1B
Complete this question by entering your answers in the tabs below Ree 1 Ree 182Ree 28Re 3 For direct materials, the materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract? OYes ONo Req 1A Req 2A
Complete this question by entering your answers in the tabs below. Req 1AReq 1B Req 2A Rq 2BReq 3 For direct labor, compute the rate and efficiency variances. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.) Labor rate variance Labor efficiency variance Req 1B Req 2B
Complete this question by entering your answers in the tabs below Req 1AReq 1B Req 2AReq 2B Req 3 In the past, the 20 technicians employed in the production of Fludex consisted of 4 senior technicians and 16 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued? OYes ONo 〈 Req 2A Req 3 〉
Complete this question by entering your answers in the tabs below Req 1AReq 1B Req 2AReq 2B Req 3 Compute the variable overhead rate and efficiency variances. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.) Variable overhead rate variance Variable overhead efficiency variance KReq 2B
Answer #1
1-a. Material price variance=Actual material used*(Standard rate for material-Actual rate for material)
Actual material used=Beginning materials inventory+Material purchased-Ending materials inventory=0+10000-3000=7000 ounce
Standard rate=$ 30 per ounce
Actual rate=Total cost of material purchase/Material purchased in ounce=287000/10000=$ 28.7 per ounce
Material price variance=7000*(30-28.7)=$9100 (F)
(Since actual rate is less than standard rate, variance is favorable)
Material quantity variance=Standard rate for materials*(Standard materials required-Actual material used)
Standard rate=$ 30 per ounce
Standard materials required=Actual units produced*Material required per unit=3400*2=6800 ounce
Actual material used=7000 ounce
Material quantity variance=30*(6800-7000)=6000 U
(Actual material used is more than the standard material required.Hence, variance is unfavorable)
1-b. Yes Since the material price variance is favorable
2-a. Labor rate variance=Actual hours worked*(Standard labor rate-Actual labor rate)
Actual hours worked=20*130=2600 hours
Standard labor rate=$14 per hour
Actual labor rate=$12 per hour
Labor rate variance=2600*(14-12)=$5200 (F)
(Actual labor rate is less than standard labor rate.Hence variance is favorable)
Labor efficiency variance=Standard labor rate*(Standard labor hours required-Actual labor hours worked)
Standard labor rate=$14 per hour
Standard labor hours required=Actual units produced*Labor hours required per unit=3400*0.5=1700 hours
Actual hours worked=2600 hours
Labor efficiency variance=14*(1700-2600)=$12600 U
(Actual hours worked is more than the standard hours required.Hence, variance is unfavorable)
2-b. No Since the Labor efficiency variance is unfavorable
3-a. Variable overhead rate variance=Actual labor hours *(Standard variable overhead rate-Actual variable overhead rate)
Actual hours worked=2600 hours
Standard variable overhead rate=$3.4 per hour
Actual variable overhead rate=Actual variable overhead cost/Actual hours worked=4700/2600=$1.81 per hour
Variable overhead rate variance=2600*(3.4-1.8)=$ 4160 F
(Actual variable overhead rate is less than standard variable overhead rate.Hence variance is favorable)
Variable overhead efficiency variance=Standard variable overhead rate*(Standard hours-Actual hours)
Standard variable overhead rate=$3.4 per hour
Standard labor hours required=Actual units produced*Labor hours required per unit=3400*0.5=1700 hours
Actual hours worked=2600 hours
Variable overhead efficiency variance=3.4*(1700-2600)=$3060 U
(Actual hours worked is more than the standard hours required.Hence, variance is unfavorable)